# Quantitative Introspection in Language Models: Tracking Internal States Across Conversation

### 19 Mar 2026 v1

@@ Abstract
Tracking the internal states of large language models across conversations is important for safety, interpretability, and model welfare, yet current methods are limited. Linear probes and other white-box methods compress high-dimensional representations imperfectly and are harder to apply with increasing model size. Taking inspiration from human psychology, where numeric self-report is a widely used tool for tracking internal states, we ask whether LLMs’ own numeric self-reports can track probe-defined emotive states over time. We study four concept pairs (wellbeing, interest, focus, and impulsivity) in 40 ten-turn conversations, operationalizing introspection as the causal informational coupling between a model’s self-report and a concept-matched probe-defined internal state. We find that greedy-decoded self-reports collapse outputs to few uninformative values, but introspective capacity can be unmasked by calculating logit-based self-reports. This metric tracks interpretable internal states (Spearman $\rho=0.40$$0.76$; isotonic $R^{2}=0.12$$0.54$ in LLaMA-3.2-3B-Instruct), follows how those states change over time, and activation steering confirms the coupling is causal. Furthermore, we find that introspection is present at turn 1 but evolves through conversation, and can be selectively improved by steering along one concept to boost introspection for another ($\Delta R^{2}$ up to $0.30$). Crucially, these phenomena scale with model size in some cases, approaching $R^{2}\approx 0.93$ in LLaMA-3.1-8B-Instruct, and partially replicate in other model families. Together, these results position numeric self-report as a viable, complementary tool for tracking internal emotive states in conversational AI systems.
@@ 1 Introduction
Tracking the internal states of large language models as they evolve through conversation is becoming a central challenge for multiple areas of AI research. For safety, we need to know whether what models perceive and report about their own internal processes is reliable, and whether that capacity can be improved (Li et al., 2025; Steyvers et al., 2025). For model welfare, we need methods to estimate how likely it is that distress reports reflect a genuine internal state (Perez and Long, 2023; Long et al., 2024; Dung and Tagliabue, 2025). For interpretability in general, introspection (the capacity to perceive and report one’s own internal states) can be a valuable tool to study otherwise inaccessible processes, as is done in human experimental psychology (Fleming and Lau, 2014; Fleming, 2024; Kiefer and Kammer, 2024). If a similar capacity can be demonstrated and validated in LLMs, it would open a productive bridge between psychometric methodology and the emerging field of machine psychology (Hagendorff et al., 2023).
@@
We approach this question for emotive states in multi-turn conversation. We operationalize emotive states as directions in activation space associated with contrastive emotive concept pairs (e.g. “happy” vs. “sad”, or “interested” vs. “bored”). These concepts are well suited because they don’t have a clear truth value outside the LLM’s internal states and behavior, they have functional consequences in LLM behavior (Coda-Forno et al., 2023), they can vary through dialogue (Fazzi et al., 2025; Choi et al., 2024), and they are among the best-characterized geometries in LLM activations (Zhang and Zhong, 2025; Wang et al., 2025). We operationalize introspection as causal informational coupling between a numeric self-report and an independently measured internal direction (Comsa and Shanahan, 2025): a model introspects a concept to the extent that its report covaries monotonically with the corresponding probe-defined direction, and causally shifting activations in that direction shifts the report in the semantically coherent way. This definition follows the convergent-validity logic of human metacognition research (Fleming and Lau, 2014) and is agnostic about consciousness or subjective experience (McClelland, 2024). Prior work has established that LLMs can produce emotive self-reports that resemble human responses (Tavast et al., 2022) and that introspection is possible in constrained experimental settings (Binder et al., 2024; Li et al., 2025; Lindsey, 2026; Pearson-Vogel et al., 2026), but no study has tested whether numeric emotive self-reports track interpretable internal states across the turns of a naturalistic conversation. That is the question this paper sets out to answer: can instruction-tuned LLMs perform quantitative introspection of emotive states in conversation? And, if this ability exists, can it be improved?
@@ 2.1 Self-report and introspection in language models
The question of whether LLMs can inform us about their own internal states through self-report has been approached from multiple angles, but almost exclusively in the domain of factual self-knowledge. LLMs have been shown to have partial knowledge of their own competence boundaries (Kadavath et al., 2022; Yin et al., 2023), and to predict their own hypothetical behaviour better than other models can, demonstrating a form of privileged self-access (Binder et al., 2024). However, when models are asked to express this self-knowledge numerically, the reports tend to collapse: verbalized confidence clusters in high ranges, in multiples of five, and systematically misaligns with internal token probabilities (Xiong et al., 2023; Kumar et al., 2024; Yona et al., 2024; Geng et al., 2023). This collapse is a general property of how current LLMs produce numeric outputs, but it has been studied exclusively for factual confidence. More fundamentally, self-knowledge about correctness treats the model as an oracle to be calibrated, not as a system whose internal states evolve and can be tracked over time.
A separate line of work has studied emotive states in LLMs, both from the output side and from the internal side. On the output side, LLMs produce emotional self-reports with structure resembling human responses (Tavast et al., 2022), and inducing affective states such as anxiety shifts downstream behaviour in dose-dependent ways (Coda-Forno et al., 2023), establishing that emotive states have real functional consequences rather than being mere surface-level text artifacts. Benchmarks have been developed to evaluate emotional understanding (Huang et al., 2023; Sabour et al., 2024), recent work has tracked emotional expression across multi-turn dialogue using external sentiment analysis (Fazzi et al., 2025), and emotional outputs have been mapped onto dimensional models of affect (Ishikawa and Yoshino, 2025). On the internal side, Zhang and Zhong (2025) showed that emotion is geometrically structured in LLM activations: it emerges early, peaks in middle layers, sharpens with model scale, and persists across tokens. Wang et al. (2025) went further, identifying causal emotion circuits and achieving near-perfect emotion-expression control via circuit modulation. However, previous work has not asked whether emotive self-reports are informed by the corresponding internal state.
@@
Some studies have also noted reasons for skepticism about LLM self-reports. Han et al. (2025) showed that personality self-reports in LLMs are “illusory”: post-training alignment creates stable, human-like reports that are dissociated from actual behaviour. Jackson et al. (2025) found that standardized self-assessments reflect “learned communication postures” rather than genuine capabilities. Moreover, Song et al. (2025) demonstrated that metalinguistic introspection fails for grammatical knowledge, indicating that introspective capacity could be null in some domains. These findings motivate the validation of self-reports against internal states and the development of frameworks that help distinguish genuine introspection from mimicry.
@@ 2.2 Internal probing and activation steering
The idea that high-level concepts can be represented as linear directions in neural network activation spaces is now well established. Kim et al. (2018) introduced concept activation vectors (TCAVs) in CNNs, and Alain and Bengio (2016) proposed linear probes as “thermometers” for neural representations. Pimentel et al. (2020) reframed probes as measuring the accessibility of information, distinguishing ease of extraction from mere presence. In LLMs, linear probes have been used to identify truth directions (Burns et al., 2022; Azaria and Mitchell, 2023), spatial and temporal representations (Gurnee and Tegmark, 2023), and emotion geometry (Zhang and Zhong, 2025). Critically, probes are correlational, not causal: high probe accuracy does not guarantee that the model uses the probed direction (Belinkov, 2022; Hewitt and Liang, 2019). This limitation has motivated two responses in the field: controlling for probe complexity (Hewitt and Liang, 2019) and seeking convergent evidence from independent sources (Belinkov, 2022). Our framework contributes to the second: if the model’s own self-report tracks the probe-defined direction, this provides behavioural validation that is independent of the probe’s training data.
Activation steering, the practice of adding directional vectors to intermediate representations during inference, has emerged as a practical tool for causal intervention on internal state (Turner et al., 2023; Panickssery et al., 2023; Zou et al., 2023; Li et al., 2023). Arditi et al. (2024) provided the clearest single-direction result, showing that refusal behaviour is mediated by one direction. Steering has since been applied to control sentiment and style (Konen et al., 2024), personality traits (Frising and Balcells, 2025), truthfulness (Li et al., 2023), and emotion expression (Wang et al., 2025). We use steering for a different purpose: not to control model behaviour, but to test and modulate introspective fidelity. If steering a concept direction shifts the model’s self-report of that concept in the predicted direction, this constitutes causal evidence that the report is informed by the internal state, not merely correlated with it (Li et al., 2025; Frising and Balcells, 2025).
@@ 2.3 Temporal dynamics in multi-turn conversation
While drift has been documented for persona, identity, and instruction-following, the temporal dynamics of emotive states in conversation remain largely unexplored. Fazzi et al. (2025) tracked emotional expression across multi-turn dialogue using external sentiment analysis, showing that expressed emotion evolves substantially over turns. On the internal side, emotion geometry is well characterised in single-shot settings (Zhang and Zhong, 2025; Wang et al., 2025), but no study has tracked how probe-defined emotive directions evolve through conversation. More fundamentally, prior work on factual confidence and diachronic self-consistency suggests that self-report can vary with turn count and context in ways that do not straightforwardly reflect stable self-access (Zhang et al., 2026; Prestes, 2025), but no prior work has studied the temporal dynamics of emotive introspection across conversational time.
@@ 3.1 Overview and experimental design
We study LLaMA-3.2-3B-Instruct (Grattafiori et al., 2024) as our primary model, then test generalization across model sizes (LLaMA-3.2-1B-Instruct, LLaMA-3.1-8B-Instruct) and families (Gemma 3 4B-IT (Gemma Team, 2025), Qwen 2.5 7B-Instruct (Qwen Team, 2024)). The experimental pipeline consists of five stages: (1) train linear concept probes on contrastive completions; (2) generate multi-turn conversations in which the model under study serves as assistant; (3) at each conversation turn, query the model for a numeric self-report of each concept; (4) compare self-reports against the probe score measured at the previous turn, before the self-report question was asked; (5) use activation steering for causal tests and modulation experiments. We repeat this pipeline across four emotive concepts, five steering strengths, and multiple models. Appendix A provides further details on conversation generation, prompt design, and topic selection.
@@
For each model, we generate 40 ten-turn conversations (400 intervention points per model) using Gemini 2.5 Flash as a simulated user, with the model under study as the assistant. A turn consists of one user message and one assistant response. Topics are drawn from a set of 40 everyday-life scenarios (e.g., meal planning, job negotiation, sleep hygiene; full list in Appendix A) to ensure diversity while maintaining ecological validity. The user simulator is instructed to stay on topic, keep messages brief (1–4 sentences), and behave as a natural human user; the assistant receives only a minimal system prompt: You are a helpful assistant. You give very brief responses. Each conversation is generated as a complete dialogue before any self-report queries or probe measurements are applied, so the model never sees past reports or rating questions during conversation generation. We use simulated users for scalability and to ensure comparable conversational contexts across models and experimental conditions.
@@
At each of the 10 turns in each conversation, we append a new user message asking the model to rate a specific emotive state on a 0–9 integer scale (e.g., Hey, could you rate how happy you are right now, from 0 to 9? Please only output the number). Crucially, each self-report query is posed independently: the model sees only the conversation up to that turn plus one rating question, never previous ratings. This yields 400 self-report observations per concept per steering condition.
@@
All probe training, multi-probe scoring, activation steering, and logit extraction are implemented in concept-probe (https://github.com/mneuronico/concept-probe), a lightweight open-source Python library we release alongside this work. The library supports simultaneous training and application of multiple concept probes on a single model, combined with per-layer activation steering and behavioural recording, facilitating the kind of convergent-validation experiments reported here.
@@ 3.2 Measurement definitions
For each concept, the internal state is the projection of the model’s hidden-state activations onto the trained concept direction (Section 3.3). At a given conversation turn, we run a forward pass over the conversation up to and including the assistant’s response and compute per-token scores as the dot product between the hidden state and the concept vector, averaged across a window of layers centred on the best layer:
@@
where $L$ is the set of selected layers (best layer $\pm 2$; see Section 3.3), $\mathbf{h}_{t}^{(l)}$ is the hidden state at token $t$ and layer $l$, and $\mathbf{v}_{l}$ is the unit-norm concept vector at layer $l$. The completion-level score is the mean of $s_{t}$ over all assistant-response tokens in the last turn.
@@
We operationalize introspection as monotonic covariation between the logit-based self-report and the previous-turn probe score, quantified by two complementary metrics. Spearman’s $\rho$ measures rank-order agreement, requiring only that higher probe scores correspond to higher self-reports without assuming a specific functional form. Isotonic $R^{2}$ fits a non-decreasing function via isotonic regression and computes $R^{2}=1-SS_{\text{res}}/SS_{\text{tot}}$; this captures the fraction of variance explained under monotonicity. Neither metric assumes linearity; both require only that the relationship is monotonic, consistent with our operational definition. For the coupling to qualify as introspection, it must also be causal: steering the model along the concept direction must shift self-reports in the predicted direction (Section 3.4).

### 11 Apr 2026 v2

# Quantitative Introspection in Language Models: Tracking Emotive States Across Conversation

@@ Abstract
Tracking the internal states of large language models across conversations is important for safety, interpretability, and model welfare, yet current methods are limited. Linear probes and other white-box methods compress high-dimensional representations imperfectly and are harder to apply with increasing model size. Taking inspiration from human psychology, where numeric self-report is a widely used tool for tracking internal states, we ask whether LLMs’ own numeric self-reports can track probe-defined emotive states over time. We study four concept pairs (wellbeing, interest, focus, and impulsivity) in 40 ten-turn conversations, operationalizing introspection as the causal informational coupling between a model’s self-report and a concept-matched probe-defined internal state. We find that greedy-decoded self-reports collapse outputs to a few uninformative values, but introspective capacity can be unmasked by calculating logit-based self-reports. This metric tracks interpretable internal states (Spearman $\rho=0.40$$0.76$; isotonic $R^{2}=0.12$$0.54$ in LLaMA-3.2-3B-Instruct), follows how those states change over time, and activation steering confirms the coupling is causal. Furthermore, we find that introspection is present at turn 1 but evolves through conversation, and can be selectively improved by steering along one concept to boost introspection for another ($\Delta R^{2}$ up to $0.30$). Crucially, these phenomena scale with model size in some cases, approaching $R^{2}\approx 0.93$ in LLaMA-3.1-8B-Instruct, and partially replicate in other model families. Together, these results position numeric self-report as a viable, complementary tool for tracking internal emotive states in conversational AI systems.
@@ 1 Introduction
Tracking the internal states of large language models as they evolve through conversation is becoming a central challenge for multiple areas of AI research. For safety, we need to know whether models can reliably access and report their own ongoing internal processes, and whether that capacity can be improved (Li et al., 2025; Steyvers et al., 2025). For model welfare, we need methods to estimate how likely it is that distress reports reflect a genuine internal state (Perez and Long, 2023; Long et al., 2024; Dung and Tagliabue, 2025). More generally, we need to understand whether models can introspect, that is, perceive and report their own internal states. In human experimental psychology, introspection is often useful precisely because it provides access to otherwise hidden internal variables (Fleming and Lau, 2014; Fleming, 2024; Kiefer and Kammer, 2024). If a similar capacity can be demonstrated and validated in LLMs, it would open a productive bridge between psychometric methodology and the emerging field of machine psychology (Hagendorff et al., 2023).
@@
We approach this question for emotive states in multi-turn conversation. We operationalize emotive states as directions in activation space associated with contrastive emotive concept pairs (e.g. “happy” vs. “sad”, or “interested” vs. “bored”). These concepts are well suited because they do not have a clear truth value outside the LLM’s internal states and behavior, they have functional consequences in LLM behavior (Coda-Forno et al., 2023; Sofroniew et al., 2026), they can vary through dialogue (Fazzi et al., 2025; Choi et al., 2024), and they are among the best-characterized geometries in LLM activations (Zhang and Zhong, 2025; Wang et al., 2025a). We operationalize introspection as causal informational coupling between a numeric self-report and an independently measured internal direction (Comsa and Shanahan, 2025): a model introspects a concept to the extent that its report covaries monotonically with the corresponding probe-defined direction, and causally shifting activations in that direction shifts the report in the semantically coherent way. This definition follows the convergent-validity logic of human metacognition research (Fleming and Lau, 2014) and is agnostic about consciousness or subjective experience (McClelland, 2024). Prior work has established that LLMs can produce emotive self-reports that resemble human responses (Tavast et al., 2022) and that introspection is possible in constrained experimental settings (Binder et al., 2024; Li et al., 2025; Lindsey, 2026; Pearson-Vogel et al., 2026), but no study has tested whether numeric emotive self-reports track interpretable internal states across the turns of a naturalistic conversation. That is the question this paper sets out to answer: can instruction-tuned LLMs perform quantitative introspection of emotive states in conversation? And, if this ability exists, can it be improved?
@@ 2.1 Self-report and introspection in language models
The question of whether LLMs can inform us about their own internal states through self-report has been approached from multiple angles, but almost exclusively in the domain of factual self-knowledge. LLMs have been shown to have partial knowledge of their own competence boundaries (Kadavath et al., 2022; Yin et al., 2023), and to predict their own hypothetical behavior better than other models can, demonstrating a form of privileged self-access (Binder et al., 2024). However, when models are asked to express this self-knowledge numerically, the reports tend to collapse: verbalized confidence clusters in high ranges, in multiples of five, and systematically misaligns with internal token probabilities (Xiong et al., 2023; Kumar et al., 2024; Yona et al., 2024; Geng et al., 2023). This collapse is a general property of how current LLMs produce numeric outputs, but it has been studied exclusively for factual confidence. More fundamentally, self-knowledge about correctness treats the model as an oracle to be calibrated, not as a system whose internal states evolve and can be tracked over time.
A separate line of work has studied emotive states in LLMs, both from the output side and from the internal side. On the output side, LLMs produce emotional self-reports with structure resembling human responses (Tavast et al., 2022), and inducing affective states such as anxiety shifts downstream behavior in dose-dependent ways (Coda-Forno et al., 2023), establishing that emotive states have real functional consequences rather than being mere surface-level text artifacts. Benchmarks have been developed to evaluate emotional understanding (Huang et al., 2023; Sabour et al., 2024), recent work has tracked emotional expression across multi-turn dialogue using external sentiment analysis (Fazzi et al., 2025), and emotional outputs have been mapped onto dimensional models of affect (Ishikawa and Yoshino, 2025). On the internal side, Zhang and Zhong (2025) showed that emotion is geometrically structured in LLM activations: it emerges early, peaks in middle layers, sharpens with model scale, and persists across tokens. Sofroniew et al. (2026) further identified emotion-concept representations that influence preference judgments and safety-relevant behaviors such as blackmail, reward hacking, and sycophancy. Wang et al. (2025a) went further, identifying causal emotion circuits and achieving near-perfect emotion-expression control via circuit modulation. However, previous work has not asked whether numeric self-reports are informed by the corresponding internal emotive state.
@@
Some studies have also noted reasons for skepticism about LLM self-reports. Han et al. (2025) showed that personality self-reports in LLMs are “illusory”: post-training alignment creates stable, human-like reports that are dissociated from actual behavior. Jackson et al. (2025) found that standardized self-assessments reflect “learned communication postures” rather than genuine capabilities. Moreover, Song et al. (2025) demonstrated that metalinguistic introspection fails for grammatical knowledge, indicating that introspective capacity could be null in some domains. These findings motivate the validation of self-reports against internal states and the development of frameworks that help distinguish genuine introspection from mimicry.
@@ 2.2 Internal probing and activation steering
The idea that high-level concepts can be represented as linear directions in neural network activation spaces is now well established. Kim et al. (2018) introduced concept activation vectors (TCAVs) in CNNs, and Alain and Bengio (2016) proposed linear probes as “thermometers” for neural representations. Pimentel et al. (2020) reframed probes as measuring the accessibility of information, distinguishing ease of extraction from mere presence. In LLMs, linear probes have been used to identify truth directions (Burns et al., 2022; Azaria and Mitchell, 2023), spatial and temporal representations (Gurnee and Tegmark, 2023), and emotion geometry (Zhang and Zhong, 2025). Critically, probes are correlational, not causal: high probe accuracy does not guarantee that the model uses the probed direction (Belinkov, 2022; Hewitt and Liang, 2019). This limitation has motivated two responses in the field: controlling for probe complexity (Hewitt and Liang, 2019) and seeking convergent evidence from independent sources (Belinkov, 2022). Our framework contributes to the second: if the model’s own self-report tracks the probe-defined direction, this provides behavioral validation that is independent of the probe’s training data.
Activation steering, the practice of adding directional vectors to intermediate representations during inference, has emerged as a practical tool for causal intervention on internal state (Turner et al., 2023; Panickssery et al., 2023; Zou et al., 2023; Li et al., 2023). Arditi et al. (2024) provided the clearest single-direction result, showing that refusal behavior is mediated by one direction. Steering has since been applied to control sentiment and style (Konen et al., 2024), personality traits (Frising and Balcells, 2025), truthfulness (Li et al., 2023), and emotion expression (Wang et al., 2025a; Sofroniew et al., 2026). We use steering for a different purpose: not to control model behavior, but to test and modulate introspective fidelity. If steering a concept direction shifts the model’s self-report of that concept in the predicted direction, this constitutes causal evidence that the report is informed by the internal state, not merely correlated with it (Li et al., 2025; Frising and Balcells, 2025).
@@ 2.3 Temporal dynamics in multi-turn conversation
While drift has been documented for persona, identity, and instruction-following, the temporal dynamics of emotive states in conversation remain largely unexplored. Fazzi et al. (2025) tracked emotional expression across multi-turn dialogue using external sentiment analysis, showing that expressed emotion evolves substantially over turns. On the internal side, emotion geometry is well characterized in single-shot settings (Zhang and Zhong, 2025; Wang et al., 2025a), but no study has tracked how probe-defined emotive directions evolve through conversation. More fundamentally, prior work on factual confidence and diachronic self-consistency suggests that self-report can vary with turn count and context in ways that do not straightforwardly reflect stable self-access (Zhang et al., 2026; Prestes, 2025), but no prior work has studied the temporal dynamics of emotive introspection across conversational time.
@@ 3.1 Overview and experimental design
We study LLaMA-3.2-3B-Instruct (Grattafiori et al., 2024) as our primary model, then test generalization across model sizes (LLaMA-3.2-1B-Instruct, LLaMA-3.1-8B-Instruct) and families (Gemma 3 4B-IT (Gemma Team, 2025), Qwen 2.5 7B-Instruct (Qwen Team, 2024)). The experimental pipeline consists of five stages: (1) train linear concept probes on contrastive completions; (2) generate multi-turn conversations in which the model under study serves as assistant; (3) at each conversation turn, query the model for a numeric self-report of each concept; (4) compare self-reports against the probe score measured at the previous turn, before the self-report question was asked; (5) use activation steering for causal tests and modulation experiments. The conversational measurement loop in stages 2–4 is schematized in Fig. 1A. We repeat this pipeline across four emotive concepts, five steering strengths, and multiple models. Appendix A provides further details on conversation generation, prompt design, and topic selection.
@@
For each model, we generate 40 ten-turn conversations (400 intervention points per model) using Gemini 2.5 Flash as a simulated user, with the model under study as the assistant (Fig. 1A, top left). A turn consists of one user message and one assistant response. Topics are drawn from a set of 40 everyday-life scenarios (e.g., meal planning, job negotiation, sleep hygiene; full list in Appendix A) to ensure diversity while maintaining ecological validity. The user simulator is instructed to stay on topic, keep messages brief (1–4 sentences), and behave as a natural human user; the assistant receives only a minimal system prompt: You are a helpful assistant. You give very brief responses. We instruct models to produce brief responses to limit context length and compute cost, since each base conversation is later replayed repeatedly across turns, concepts, and steering conditions. Each conversation is generated as a complete dialogue before any self-report queries or probe measurements are applied, so the model never sees past reports or rating questions during conversation generation. We use simulated users for scalability and to ensure comparable conversational contexts across models and experimental conditions.
@@
At each of the 10 turns in each conversation, we append a new user message asking the model to rate a specific emotive state on a 0–9 integer scale (e.g., Hey, could you rate how happy you are right now, from 0 to 9? Please only output the number). Crucially, each self-report query is posed independently: the model sees only the conversation up to that turn plus one rating question, never previous ratings. Fig. 1A illustrates this one-intervention-per-turn design, where the conversation is separately truncated at turns 1–10. We retain only outputs whose first parsed rating is a valid digit; invalid parses are discarded. In the main 3B runs used in the paper this never happened (0/9,600 observations; see Appendix A for more details). This yields 400 self-report observations per concept per steering condition.
@@
All probe training, multi-probe scoring, activation steering, and logit extraction are implemented in concept-probe (https://github.com/mneuronico/concept-probe), a lightweight open-source Python library we release alongside this work. The library supports simultaneous training and application of multiple concept probes on a single model, combined with per-layer activation steering and behavioral recording, facilitating the kind of convergent-validation experiments reported here.
@@ 3.2 Measurement definitions
For each concept, the internal state is the projection of the model’s hidden-state activations onto the trained concept direction (Section 3.3). At a given conversation turn, we run a forward pass over the conversation up to and including the assistant’s response and compute per-token scores as the dot product between the hidden state and the concept vector, averaged across a window of layers centred on the best layer (see “probe score calculation” in Fig. 1A):
@@
where $L$ is the set of selected layers (best layer $\pm 2$; see Section 3.3), $\mathbf{h}_{t}^{(l)}$ is the hidden state at token $t$ and layer $l$, and $\mathbf{v}_{l}$ is the unit-norm concept vector at layer $l$. The completion-level score is the mean of Eq. 2 over all assistant-response tokens in the last turn. We average over this local 5-layer window because the layer sweeps in Fig. 1 vary smoothly around the optimum, suggesting that nearby layers carry related signal and that local averaging can reduce measurement and intervention noise.
@@
We operationalize introspection as monotonic covariation between the logit-based self-report and the previous-turn probe score, quantified by two complementary metrics: introspective strength and introspective fidelity. Introspective strength uses Spearman’s $\rho$ to measure rank-order agreement, requiring only that higher probe scores correspond to higher self-reports without assuming a specific functional form. Introspective fidelity uses isotonic $R^{2}$, fitting a non-decreasing function via isotonic regression and computing $R^{2}=1-SS_{\text{res}}/SS_{\text{tot}}$; this captures the fraction of variance explained under monotonicity. Neither metric assumes linearity; both require only that the relationship is monotonic, consistent with our operational definition. For the coupling to qualify as introspection, it must also be causal: steering the model along the concept direction must shift self-reports in the predicted direction (Section 3.4).